# VasoMIM: Vascular Anatomy-Aware Masked Image Modeling for Vessel Segmentation

### v1

@@ Introduction
Over the past decade, numerous vessel segmentation algorithms have been proposed (Ronneberger, Fischer, and Brox 2015; Huang et al. 2024; Wu et al. 2025), helping to alleviate radiologists’ workload. However, training high-performance models requires large-scale datasets of X-ray angiograms with pixel-level labels, and producing such annotations remains labor-intensive, time-consuming, and dependent on specialized domain knowledge (Esteva et al. 2019). Self-supervised learning (SSL) provides an attractive alternative by learning generalizable representations from vast unlabeled data, thereby boosting downstream performance (Gui et al. 2024). In particular, masked image modeling (MIM) (He et al. 2022; Fu et al. 2025; Zhuang et al. 2025b; Tang et al. 2025) has achieved remarkable success in natural and medical image analysis by training models to reconstruct masked image patches, as shown in Fig. 1 (a).
@@
  • Extensive experiments demonstrate the benefits of integrating anatomical knowledge into the MIM framework. Meanwhile, VasoMIM consistently outperforms state-of-the-art SSL alternatives on vessel segmentation tasks [^1] VasoMIM also exhibits superior performance on image denoising tasks (see Supplementary Material)..
@@
Self-supervised learning (SSL) (Chen et al. 2020; Chen, Xie, and He 2021; Caron et al. 2021; He et al. 2022; Wang et al. 2025) offers a practical solution to mitigate annotation scarcity in medical imaging. Existing approaches can be categorized into two main paradigms: contrastive learning (CL)-based and masked image modeling (MIM)-based. CL-based methods aim to pull positive pairs together while pushing negative pairs apart in feature space, such as C22(Zhou et al. 2020), MICLe (Azizi et al. 2021), VoCo (Wu, Zhuang, and Chen 2024), RAD-DINO (Pérez-Garc´ıa et al. 2025), etc. However, because these methods focus on image-level representations, they may perform poorly on dense prediction tasks like segmentation (Chaitanya et al. 2020). MIM-based approaches train models to reconstruct masked patches, yielding finer-grained representations (Yuan et al. 2023). The choice of training objective is critical. Most methods use a pixel-level reconstruction loss (Kang et al. 2024; Li et al. 2024; Xu et al. 2025), while some incorporate a contrastive loss to improve multi-view alignment in 3D medical images (Zhuang et al. 2025a, b). However, these objectives do not ensure semantic consistency in the reconstructed images, often resulting in weak vascular representations. Our work addresses this limitation by introducing a semantic-level anatomical consistency loss.
@@
The design of the masking strategy is crucial in MIM given the high redundancy of images (He et al. 2022). Most methods (Xie et al. 2024; Kang et al. 2024) adopt random masking at high ratios (e.g., $75\%$). To create more challenging pretext tasks, many carefully designed masking strategies have been proposed. AMT (Liu, Gui, and Luo 2023) masked patches based on attention maps. SemMAE (Li et al. 2022a) learned semantic parts of images first and used part segmentation results to guide patch masking. HAP (Yuan et al. 2023) exploited human-structure priors to guide the masking process for human-centric perception pre-training. Methods such as HPM (Wang et al. 2025), AnatoMask (Li et al. 2024) and AHM (Xu et al. 2025) identified patches that are difficult to reconstruct through a loss predictor, a self-distillation framework and a policy network, respectively, and preferentially masked these hard patches. However, none of these approaches can explicitly incorporate vascular anatomy into the masking process, which is essential for guiding models to focus on vascular regions.
@@
Traditional vessel segmentation methods, such as Frangi filter (Frangi et al. 1998), active contours (Taghizadeh Dehkordi et al. 2014) and graph cuts (Wang et al. 2020), rely on handcrafted features that generalize poorly across the wide variability of X-ray angiograms. The emergence of deep learning has significantly advanced the field. Early CNN-based architectures like U-Net (Ronneberger, Fischer, and Brox 2015) and its variants (Zhou et al. 2019; Li et al. 2020) capture hierarchical features but remain constrained by their local receptive fields. More recent transformer- and Mamba-based models (Chen et al. 2024; Hatamizadeh et al. 2021; Ruan, Li, and Xiang 2024; Wang et al. 2024) address this limitation by integrating global context, thereby improving delineation of complex vascular structures. However, these deep models are highly data-hungry, and the scarcity of labeled angiograms continues to limit their gains (Zhou et al. 2021). In this work, we leverage MIM to learn generalizable representations from large-scale unlabeled data, thereby improving the performance of existing segmentors.
@@
Although the threshold $\alpha$ ($92$ by default) influences the quality of the extracted vascular anatomy, ablations provided in Supplementary Material demonstrate that VasoMIM is robust to variations in $\alpha$.
@@
This strategy ensures the model initially encounters a balanced mix of vessel and background patches, and then progressively focuses on vascular regions, thereby improving its ability to learn vascular representations.
@@
Pre-Training ARCADE (Popov et al. 2024) $2,000$ -
CADICA (Jiménez-Partinen et al. 2024) $6,594$ -
Stenosis (Danilov et al. 2021) $7,492$ -
SYNTAX (Mahmoudi et al. 2025) $2,943$ -
XCAD (Ma et al. 2021) $1,621$ -
@@
Vessel Segmentation ARCADE (Popov et al. 2024) $200$ $300$
XCAV (Wu et al. 2025) $175$ $46$
CAXF (Li et al. 2020) $337$ $201$
@@ Results
Table 1 summarizes the datasets used for both pre-training and vessel segmentation. More detailed descriptions can be found in Supplementary Material.
@@
Vessel Segmentation. VasoMIM is evaluated on three benchmarks, including one in-domain dataset, ARCADE, and two out-of-domain datasets, CAXF (Li et al. 2020) and XCAV (Wu et al. 2025). Note that the images of ARCADE used for vessel segmentation are not included in the pre-training set.
@@
Frangi Filter (Frangi et al. 1998) [MICCAI’98] $41.30$ $40.91$ $64.01$ $65.73$ $58.46$ $57.15$
@@
[…]
Table 2: Comparison of state-of-the-art methods on ARCADE, CAXF and XCAV. All methods are reimplemented using their official codebases. The best and second-best results are highlighted in red and yellow, respectively. Results are reported as ${\rm mean}\pm{\rm std}$” over three random seeds, except for Frangi filter. indicates that the model is specialized in medical imaging.
@@
Vessel Segmentation. We adopt U-Net (Ronneberger, Fischer, and Brox 2015) as the segmentation decoder and fine-tune it end-to-end on ARCADE, CAXF and XCAV for $500$ epochs each, using input images resized to $224\times 224$. Optimization is performed using AdamW with an initial learning rate of $1e^{-4}$ and a weight decay of $0.05$. A cosine-annealing schedule with TT_max = $500$ is employed. All experiments are conducted on an NVIDIA ${\rm A}6000$ GPU.
Evaluation Metrics. Dice similarity coefficient (DSC) and centerlineDice (clDice) (Shit et al. 2021) are adopted to evaluate the performance of vessel segmentation. Compared to DSC, clDice better captures topological correctness by measuring overlap between predicted and ground-truth vascular centerlines.
@@
We compare VasoMIM against both CL-based and MIM-based SSL methods. All baselines are pre-trained on the angiogram dataset in Table 1 using their official implementations. Each model is fine-tuned with three random seeds, and we report all metrics as ${\rm mean}\pm{\rm std}$”.
@@
Guidance $\mathcal{L}_{\rm cons.}$ ARCADE CAXF
DSC (%) clDice (%) DSC (%) clDice (%)
- - $68.00$ $69.78$ $83.15$ $86.41$
- $68.45$ $\uparrow 0.45$ $70.18$ $\uparrow 0.40$ $84.03$ $\uparrow 0.88$ $87.71$ $\uparrow 1.30$
- $68.30$ $\uparrow 0.30$ $69.94$ $\uparrow 0.16$ $83.96$ $\uparrow 0.81$ $87.61$ $\uparrow 1.20$
$\mathbf{68.85}$ $\mathbf{\uparrow 0.85}$ $\mathbf{70.56}$ $\mathbf{\uparrow 0.78}$ $\mathbf{84.49}$ $\mathbf{\uparrow 1.34}$ $\mathbf{88.33}$ $\mathbf{\uparrow 1.92}$
Table 3: Ablation study of the proposed anatomy-guided masking strategy and anatomical consistency loss.
@@
Qualitative Results. Several representative cases from the three datasets are presented in Fig. 3. VasoMIM produces precise vessel segmentations, even for very thin or blurred vessels. Due to space limitations, additional qualitative results are provided in Supplementary Material.
Ablation Study
Ablation studies are conducted on ARCADE and CAXF with default settings highlighted in orange. Each model is fine-tuned with three random seeds. Table 3 shows the results of sequentially adding each proposed component to the vanilla baseline, i.e., MAE ($\gamma=0.5$).
Effectiveness of Anatomy-Guided Masking Strategy. We observe that simply adding the anatomy-guided masking strategy yields a clear performance boost ($0.30\%$ DSC on ARCADE and $0.81\%$ DSC on CAXF).
@@
Case Guidance Randomness $\beta_{0}$ $\beta_{E}$ DSC (%)
ARCADE CAXF
Random weak strong 0 0 $68.45$ $84.03$
Anatomy-Guided $\downarrow$ $\downarrow$ 0 $0.5$ $\mathbf{68.85}$ $\mathbf{\uparrow 0.40}$ $\mathbf{84.49}$ $\mathbf{\uparrow 0.46}$
0 11 $68.52$ $\uparrow 0.07$ $84.24$ $\uparrow 0.21$
strong weak 11 11 $65.36$ $\downarrow 3.09$ $81.17$ $\downarrow 2.86$
Table 5: Effects of different masking strategies. As $\beta_{E}$ increases, the anatomical guidance becomes stronger while the randomness decreases.
To better understand this improvement, we measure the proportion of vessel-containing patches in the masked patches during the pre-training process. As shown in Fig. 4 (a), our strategy masks far more patches containing vessels, whereas the baseline masks a nearly fixed (small) proportion of vessel-containing patches regardless of image content. We further visualize an example in Fig. 4 (b), where the masking ratio of each patch over pre-training is colored. Our strategy clearly favors masking patches rich in vascular anatomy, whereas the baseline shows no such preference.
Role of Anatomical Consistency Loss. Adding the anatomical consistency loss $\mathcal{L}_{\rm cons.}$ to the baseline also produces a notable gain (e.g., $83.15\%\rightarrow 84.03\%$ DSC on CAXF). This improvement suggests that $\mathcal{L}_{\rm cons.}$ enables the model to learn more discriminative vascular representations. To further verify the impact of $\mathcal{L}_{\rm cons.}$, we first split patches into vessel‑containing and background-only groups, then use UMAP (Healy and McInnes 2024) to project their representations from the model pre-trained w/ and w/o $\mathcal{L}_{\rm cons.}$ into a low-dimensional space, and evaluate how well these representations cluster. As shown in Table 4, the model pre-trained w/ $\mathcal{L}_{\rm cons.}$ achieves much higher SS and CHI, and a far lower DBI than the model pre-trained w/o $\mathcal{L}_{\rm cons.}$, indicating more compact and well-separated clusters of vascular representations.
Case Manner $\beta_{0}$ $\beta_{E}$ DSC (%)
ARCADE CAXF
Random - 0 0 $68.45$ $84.03$
Anatomy-Guided Weak-to-Strong 0 $0.5$ $\mathbf{68.85}$ $\mathbf{\uparrow 0.40}$ $\mathbf{84.49}$ $\mathbf{\uparrow 0.46}$
Strong-to-Weak $0.5$ 0 $67.81$ $\downarrow 0.64$ $83.41$ $\downarrow 0.62$
Table 6: Comparison of weak-to-strong and strong-to-weak anatomical guidance.
@@
In-Depth Analysis
We next present a detailed analysis of the proposed VasoMIM. Default settings are highlighted in orange. All results are averaged over three random seeds.
@@
Different Masking Strategies. We evaluate several masking strategies, with results summarized in Table 5. Different strategies are constructed by varying $\beta_{0}$ and $\beta_{E}$. For instance, setting $\beta_{0}=\beta_{E}=0$ yields a fully random masking strategy, whereas $\beta_{0}=\beta_{E}=1$ indicates that all $\gamma N$ patches are masked purely based on anatomical guidance. Interestingly, simply increasing the degree of anatomical guidance does not consistently yield better performance. Evidently, retaining a certain degree of randomness in the masking process is essential. Specifically, the configuration $\beta_{0}=0$, $\beta_{E}=0.5$ achieves the best results, striking a balance between stronger anatomical guidance ($\beta_{0}=\beta_{E}=1$) and greater randomness ($\beta_{0}=\beta_{E}=0$). This outcome is quite intuitive. Aggressively masking patches with the highest anatomical relevance leads to most vessel-containing patches being masked. In such cases, the model is forced to reconstruct vascular anatomy solely from background-only patches with little to no semantic cues.
Figure 7: Fine-tuning models pre-trained under different epochs on CAXF. GFLOPs are evaluated on an NVIDIA $\rm A6000$} GPU, using a single $224\times 224$ RGB image (with masking) as the input.
@@
Scratch [MICCAI’15] $38.09$ $50.67$ $58.27$
MAE [CVPR’22] $56.58$ $63.04$ $68.17$
VasoMIM [Ours] $\mathbf{56.93}$ $\mathbf{63.64}$ $\mathbf{68.85}$
@@
Pre-Training and Data Efficiency. We find that VasoMIM’s advantages become even more pronounced under resource-limited scenarios, either with fewer pre-training epochs or with less annotated data for fine-tuning. First, we fine-tune MAE and VasoMIM pre-trained for $100$, $300$, $500$ and $800$ epochs on CAXF. As shown in Fig. 7, VasoMIM consistently outperforms MAE at every pre-training duration. Notably, after only $300$ epochs of pre-training ($\sim 6$ hours), VasoMIM achieves a DSC of $83.75\%$ on CAXF, slightly above MAE’s $83.53\%$, which requires $800$ epochs ($\sim 12$ hours). Even when we extend MAE’s pre-training to $1,600$ epochs ($\sim 24$ hours) for a further performance boost, VasoMIM pre-trained for only 500 epochs ($\sim 10$ hours) still performs better. Next, to test data efficiency, we fine-tune models on ARCADE using only $25\%$, $50\%$, or $100\%$ of the training labels. Table 7 shows that VasoMIM outperforms MAE in all scenarios.

### v2

@@ Introduction
Over the past decade, numerous vessel segmentation algorithms have been proposed (Ronneberger et al. 2015; Huang et al. 2024; Wu et al. 2025), helping to alleviate radiologists’ workload. However, training high-performance models requires large-scale datasets of X-ray angiograms with pixel-level labels, and producing such annotations remains labor-intensive, time-consuming, and dependent on specialized domain knowledge (Esteva et al. 2019). Self-supervised learning (SSL) provides an attractive alternative by learning generalizable representations from vast unlabeled data, thereby boosting downstream performance (Gui et al. 2024). In particular, masked image modeling (MIM) (He et al. 2022; Fu et al. 2025; Zhuang et al. 2025b; Tang et al. 2025) has achieved remarkable success in natural and medical image analysis by training models to reconstruct masked image patches, as shown in Fig. 1 (a).
@@
  • Extensive experiments demonstrate the benefits of integrating anatomical knowledge into the MIM framework. Meanwhile, VasoMIM consistently outperforms state-of-the-art SSL alternatives on vessel segmentation tasks.
@@
Self-supervised learning (SSL) (Chen et al. 2020; Chen, Xie, and He 2021; Caron et al. 2021; He et al. 2022; Wang et al. 2025) offers a practical solution to mitigate annotation scarcity in medical imaging. Existing approaches can be categorized into two main paradigms: contrastive learning (CL)-based and masked image modeling (MIM)-based. CL-based methods aim to pull positive pairs together while pushing negative pairs apart in feature space, such as C22(Zhou et al. 2020), MICLe (Azizi et al. 2021), VoCo (Wu, Zhuang, and Chen 2024), RAD-DINO (Pérez-García et al. 2025), etc. However, because these methods focus on image-level representations, they may perform poorly on dense prediction tasks like segmentation (Chaitanya et al. 2020). MIM-based approaches train models to reconstruct masked patches, yielding finer-grained representations (Yuan et al. 2023). The choice of training objective is critical. Most methods use a pixel-level reconstruction loss (Kang et al. 2024; Li et al. 2024; Xu et al. 2025), while some incorporate a contrastive loss to improve multi-view alignment in 3D medical images (Zhuang et al. 2025a, b). However, these objectives do not ensure semantic consistency in the reconstructed images, often resulting in weak vascular representations. Our work addresses this limitation by introducing a semantic-level anatomical consistency loss.
@@
The design of the masking strategy is crucial in MIM given the high redundancy of images (He et al. 2022). Most methods (Xie et al. 2024; Kang et al. 2024) adopt random masking at high ratios (e.g., $75\%$). To create more challenging pretext tasks, many carefully designed masking strategies have been proposed. AMT (Liu, Gui, and Luo 2023) masked patches based on attention maps. SemMAE (Li et al. 2022a) learned semantic parts of images first and used part segmentation results to guide patch masking. HAP (Yuan et al. 2023) exploited human-structure priors to guide the masking process for human-centric perception pre-training. Methods such as HPM (Wang et al. 2025), AnatoMask (Li et al. 2024) and, AHM (Xu et al. 2025) identified patches that are difficult to reconstruct through a loss predictor, a self-distillation framework and a policy network, respectively, and preferentially masked these hard patches. However, none of these approaches can explicitly incorporate vascular anatomy into the masking process, which is essential for guiding models to focus on vascular regions.
@@
Traditional vessel segmentation methods, such as Frangi filter (Frangi et al. 1998), active contours (Taghizadeh Dehkordi et al. 2014), and graph cuts (Wang et al. 2020), rely on handcrafted features that generalize poorly across the wide variability of X-ray angiograms. The emergence of deep learning has significantly advanced the field. Early CNN-based architectures like U-Net (Ronneberger et al. 2015) and its variants (Zhou et al. 2019; Li et al. 2020) capture hierarchical features but remain constrained by their local receptive fields. More recent transformer- and Mamba-based models (Chen et al. 2024; Hatamizadeh et al. 2021; Ruan, Li, and Xiang 2024; Wang et al. 2024) address this limitation by integrating global context, thereby improving the delineation of complex vascular structures. However, these deep models are highly data-hungry, and the scarcity of labeled angiograms continues to limit their gains (Zhou et al. 2021). This work leverages MIM to learn generalizable representations from large-scale unlabeled data, significantly boosting the performance of segmentors. We use U-Net as a representative case study to demonstrate the broad applicability of our method.
@@
Pre-Training ARCADE $2,000$ -
CADICA $6,594$ -
Stenosis $7,492$ -
SYNTAX $2,943$ -
XCAD $1,621$ -
@@
Vessel Segmentation ARCADE $200$ $300$
XCAV $175$ $46$
CAXF $337$ $201$
@@ Results
Table 1 summarizes the datasets used for both pre-training and vessel segmentation.
@@
Vessel Segmentation. VasoMIM is evaluated on three benchmarks: one in-domain dataset, ARCADE, and two out-of-domain datasets, CAXF (Li et al. 2020) and XCAV (Wu et al. 2025). Note that the images of ARCADE used for vessel segmentation are not included in the pre-training set.
@@
Frangi Filter (Frangi et al. 1998) $41.30$ $40.91$ $64.01$ $65.73$ $58.46$ $57.15$
@@
VasoMIM $\bm{68.85}$$\pm 0.47$ $\bm{70.56}$$\pm 0.36$ $\bm{84.49}$$\pm 0.17$ $\bm{88.33}$$\pm 0.09$ $\bm{77.52}$$\pm 0.26$ $\bm{74.18}$$\pm 0.34$
Table 2: Comparison of state-of-the-art methods on ARCADE, CAXF and XCAV. All methods are reimplemented using their official codebases. The best results are highlighted in bold and the second-best results are underlined. Results are reported as ${\rm mean}\pm{\rm std}$” over three random seeds, except for Frangi filter. indicates that the model is specialized in medical imaging.
@@
Vessel Segmentation. We adopt U-Net (Ronneberger et al. 2015) as the segmentation decoder and fine-tune it end-to-end on ARCADE, CAXF and XCAV for $500$ epochs each, using input images resized to $224\times 224$. Optimization is performed using AdamW with an initial learning rate of $1e^{-4}$ and a weight decay of $0.05$. A cosine-annealing schedule with TT_max = $500$ is employed. All experiments are conducted on an NVIDIA ${\rm A}6000$ GPU.
Evaluation Metrics. Dice similarity coefficient (DSC) and centerlineDice (clDice) (Shit et al. 2021) are adopted. Compared to DSC, clDice better captures topological correctness by measuring overlap between predicted and ground-truth vascular centerlines.
@@
All baselines are pre-trained on the angiogram dataset in Table 1 using their official implementations. Each model is fine-tuned with three random seeds, and we report all metrics as ${\rm mean}\pm{\rm std}$”.
@@
Guidance $\mathcal{L}_{\rm cons.}$ ARCADE CAXF
- - $68.00$ $83.15$
- $68.45$ $84.03$
- $68.30$ $83.96$
$\mathbf{68.85}$ $\mathbf{84.49}$
Table 3: Ablation study of the proposed anatomy-guided masking strategy and anatomical consistency loss. DSC is reported in this table.
@@
Qualitative Results. Several cases from the three datasets are presented in Fig. 4. VasoMIM produces precise vessel segmentations, even for very thin or blurred vessels.
Ablation Study
Ablation studies are conducted on ARCADE and CAXF with default settings highlighted in gray. Each model is fine-tuned with three random seeds. Table 3 shows the results of sequentially adding each proposed component to the vanilla baseline, i.e., MAE ($\gamma=0.5$).
Effectiveness of Anatomy-Guided Masking Strategy. Simply using the anatomy-guided masking strategy yields a clear performance boost ($0.30\%$ DSC on ARCADE and $0.81\%$ DSC on CAXF). To better understand this improvement, we measure the proportion of vessel-containing patches in the masked patches during the pre-training process. As shown in Fig. 3(a), our strategy masks a significantly larger number of vessel-containing patches, whereas the baseline masks a small and relatively constant proportion. We further visualize an example in Fig. 3 (b), where the masking ratio of each patch over pre-training is colored. Our strategy clearly favors masking patches rich in vascular anatomy, whereas the baseline shows no such preference.
Role of Anatomical Consistency Loss. Adopting the anatomical consistency loss $\mathcal{L}_{\rm cons.}$ to the baseline also produces a notable gain (e.g., $0.88\%$ DSC on CAXF).
@@
This suggests that $\mathcal{L}_{\rm cons.}$ enables the model to learn more discriminative vascular representations. To further verify the impact of $\mathcal{L}_{\rm cons.}$, we first split patches into vessel‑containing and background-only groups, then use UMAP (Healy and McInnes 2024) to project their representations from the model pre-trained w/ and w/o $\mathcal{L}_{\rm cons.}$ into a low-dimensional space, and evaluate how well these representations cluster. As shown in Table 4, the model pre-trained w/ $\mathcal{L}_{\rm cons.}$ achieves much higher SS and CHI, and a far lower DBI than the model pre-trained w/o $\mathcal{L}_{\rm cons.}$, indicating more compact and well-separated clusters of vascular representations.
In-Depth Analysis
We next present a detailed analysis of the proposed VasoMIM. Default settings are highlighted in gray. All results are averaged over three random seeds.
Case $\beta_{0}$ $\beta_{E}$ ARCADE CAXF
Random 0 0 $68.45$ $84.03$
Weak-to-Strong 0 $0.5$ $\mathbf{68.85}$ $\mathbf{84.49}$
0 11 $68.52$ $84.24$
11 11 $65.36$ $81.17$
Strong-to-Weak $0.5$ 0 $67.81$ $83.41$
Table 5: Effects of different masking strategies. DSC is reported in this table.
@@
Different Masking Strategies. We evaluate several masking strategies, with results summarized in Table 5. Simply increasing the degree of anatomical guidance does not consistently yield better performance, suggesting that a degree of randomness in the masking process is essential. Specifically, the configuration $\beta_{0}=0$, $\beta_{E}=0.5$ achieves the best results, striking a balance between stronger anatomical guidance ($\beta_{0}=\beta_{E}=1$) and greater randomness ($\beta_{0}=\beta_{E}=0$). This outcome is quite intuitive. Aggressively masking patches with the highest anatomical relevance leads to most vessel-containing patches being masked. In such cases, the model is forced to reconstruct vascular anatomy solely from background-only patches with little to no semantic cues.
Figure 7: Fine-tuning on CAXF with models pre-trained for various epochs. GFLOPs are measured on an NVIDIA $\rm A6000$} GPU using a single $224\times 224$ masked RGB image.
@@
Scratch $38.09$ $50.67$ $58.27$
MAE $56.58$ $63.04$ $68.17$
VasoMIM $\mathbf{56.93}$ $\mathbf{63.64}$ $\mathbf{68.85}$
@@
Pre-Training and Data Efficiency. VasoMIM’s advantages become even more pronounced under resource-limited scenarios. First, we fine-tune MAE and VasoMIM pre-trained for $100$, $300$, $500$, and $800$ epochs on CAXF. As shown in Fig. 7, VasoMIM consistently outperforms MAE at every pre-training duration. Notably, after only $300$ epochs of pre-training ($\sim 6$ hours), VasoMIM achieves a DSC of $83.75\%$ on CAXF, slightly above MAE’s $83.53\%$, which requires $800$ epochs ($\sim 12$ hours). Even when we extend MAE’s pre-training to $1,600$ epochs ($\sim 24$ hours) for a further performance boost, VasoMIM pre-trained for only 500 epochs ($\sim 10$ hours) still performs better. Next, to test data efficiency, we fine-tune models on ARCADE using only $25\%$, $50\%$, or $100\%$ of the training labels. Table 6 shows that VasoMIM outperforms MAE in all scenarios.